If 5 men can paint 26 houses within 3 days, how many houses can 15 men paint in 150 days
Answer by richwmiller(17219)

men times days equals work days
5*3=15 work days.
15/26 days per house
There is no way 5 guys will paint 26 houses in 3 days
that is ( 8 2/3) houses a day
26/3/5= 26/15 =1.73 houses a day per man
150/1.73=86.54 per man
15*86.4=1298 houses